c) The pore si e in apo yacrylan ide gel can be adjusted by varyingth amount ofcross linker. The %Cis the weight/ weight of total cross-linker weight in the sum of monomers and cross-linker weights. What is the %C in a bottle that has a label: 30 Acrylamide/ Bissolution. 29:1. Show all calculations ofyour answer. (2 points)
Answer #1

For 29:1 acryamide is to bis ratio

C%= 1/30 × 100 %

= 3.3.%

Weigth = 1

Weight of total cross linkers weight in sum of monomers and cross linkers weight = 29+1 =30